# Adversarial Poisoning Attack on Quantum Machine Learning Models

### v2

@@ V-A Setup
V-B Analysis
Angle vs. Amplitude Encoding: To evaluate how effectively the proposed ESS (Section IV-A) can be used to measure the performance of state preparation circuits on any given dataset, we compare the intra-class state similarity for angle and amplitude encoding under varying noise levels ($p$: error probability) on a reduced MNIST-10 dataset with a latent dimension of 16. Considering that amplitude encoding can encode $2^{n}$ features in $n$ qubits, for fairness, we used a 4-qubit circuit for angle encoding, with each qubit having four rotation gates ($R_{Z}$, $R_{X}$, $R_{Z}$, $R_{X}$), to provide a better comparison against amplitude encoding. Fig. 4 shows the performance comparison of both encoding techniques. The results are interesting, as amplitude encoding performs worse under noiseless conditions. However, it is evident that amplitude encoding outperforms angle encoding when encoding larger datasets, likely due to the loss of information caused by applying a high number of sequential rotation gates in angle encoding which is further amplified by noise. Conversely, angle encoding performs well for smaller datasets, as demonstrated in Fig. 3, where two features are encoded per qubit.
Label-flip vs. Bi-level: Even though QUID is primarily a label-flipping (LF) technique, we compare its performance against random bi-level poisoning. In this method, apart from flipping labels, we also replace samples of the poisoned data ($\mathcal{D}_{p}$) i.e. $\{x_{i}\mid x_{i}\in\mathcal{D}_{p}\subset\mathcal{D}_{tr}\}$ ($\varepsilon=0.5$), with random features and calculate the optimal poisoned labels for these random features. Table III shows the performance comparison of label-flipping versus bi-level poisoning when training a 4-qubit QNN on various datasets under noise with error rate $p=0.1$. Our experiments indicate that, in most cases, adding randomly generated poisoned samples (PS) actually leads to better performance (i.e. less accuracy & loss degradation) compared to the scenario of using only LF with QUID. For the remaining analysis, we use only the label-flipping technique.
TABLE IV: Test accuracies for QUID when trained on noiseless and noisy environment on different datasets. QUID appears to perform better under noise conditions.
@@ V-B Analysis
TABLE V: Test accuracies of QUID with poison ratio $\varepsilon=0.5$ on real hardware noise model.
@@
Noiseless vs. Noisy Simulations: We compare the performance of QUID when training a QNN under both noiseless and noisy conditions. Table. IV presents the performance comparison of QUID across various poison ratios, denoted as $\varepsilon$. From the figure, it is evident that for the same poison ratio, QUID demonstrates significantly better performance under noisy conditions compared to noiseless conditions. For example, at a poison ratio of $\varepsilon=0.5$, QUID reduces performance by $78\%$ under noisy conditions compared to 53% under noiseless conditions on Fashion-4 dataset. This result highlights the robustness of QUID against noise.
Real Hardware Noise: Given the limited availability of real quantum devices, we directly integrated noise models from actual quantum hardware. Specifically, we employed the noise models of IBM_Kyiv and IBM_Brisbane, and incorporated them into the qiskit.aer device to accurately simulate the noise characteristics of these devices. Table V lists the performance of QUID and random label-flipping on real hardware noises. We observe that even in this scenario, QUID is able to significantly degrade the performance of the baseline model by $\approx 73\%$ on IBM_Brisbane’s noise and $\approx 72\%$ on IBM_Kyiv’s noise.
TABLE VI: Comparison of test accuracy after training QNNs with various PQC architectures on different poison ratios ($\varepsilon$) under noise with error rate $p=0.05$. Cplx. refers to the complexity of the PQC in terms of circuit gate count and depth.
@@
Poison Ratio ($\boldsymbol{\varepsilon}$): We also evaluate the performance of QUID using different poison ratios, which helps determine the threshold beyond which the model’s performance decreases drastically. All training in this evaluation is conducted under noise ($p=0.05$) using the SPSA gradient calculation method. Table VI lists the test accuracies of various QNNs trained across different datasets and poison ratios ($\varepsilon$). As expected, we observe from the table that baseline models of more complex PQCs like PQC-6 perform worse compared to PQC-8 under hardware noise, while the circuit with the least gate count and depth significantly outperforms both. Consequently, we selected PQC-1 for noisy training on the other datasets. Another important observation is that with a small poison ratio ($\varepsilon\leq 0.1$), there appears to be minimal performance degradation. However, for $\varepsilon>0.5$, the performance suffers diminishing returns, as the trained model on QUID achieves a test accuracy less than $20\%$, worse than random prediction ($25\%$ for 4-class dataset), rendering the model completely unusable for the victim. In all cases, the proposed QUID technique performs significantly better than the random label-flipping technique.
TABLE VII: Test accuracies for QUID when trained w/ and w/o SS-DPA [28], a SOTA defense technique against label-flipping attacks.
@@
Dataset Baseline NoDef. SS-DPA NoDef. SS-DPA
@@
Defense Evaluation: Although label-flipping attacks are relatively easy to defend against in the classical domain using data sanitization techniques [15, 29], implementing such defenses in the current quantum cloud environment poses significant challenges, as the victim cannot intervene once the quantum circuit and training data ($\mathcal{D}_{tr}$) are submitted for training. Therefore, we evaluate the effectiveness of QUID against a practical training-based defense technique, SS-DPA [28], which is specifically designed to mitigate label-flipping attacks. SS-DPA leverages subset aggregation and ensemble learning together with semi-supervised learning. Table VII illustrates the performance of QUID with and without SS-DPA, trained in a noiseless setting. The hyperparameter $k$ in SS-DPA is set to 3 due to the smaller dataset sizes. While the application of SS-DPA improves performance by a small amount in most cases, it is not very effective for large $\varepsilon$. Moreover, it is extremely expensive in terms of computation overhead, as it requires the training and retraining of multiple QNNs. Future research should focus on the development of effective and efficient defense techniques tailored specifically to QMLs.
VI Discussions
@@
Scalability: To demonstrate the effectiveness of QUID for larger QNNs trained on larger datasets, we designed an 8-qubit QNN with two layers of PQC-1 [23]. We trained this model on the MNIST-10 dataset, reduced to a latent dimension of $d=16$, using 2000 samples with a 70:30 train-test split. Fig. 5 shows the test loss and test accuracy of the victim model when trained on a poisoned dataset with a poison ratio $\varepsilon=0.5$, using both random label flipping and QUID’s label poisoning technique. The results clearly illustrate that QUID outperforms random label flipping techniques, even for in this scenario. Specifically, QUID results in additional $\approx 41\%$ accuracy degradation and incurs $\approx 0.4$ higher test loss compared to random flipping, further validating its effectiveness.
@@
Figure 5: Test performance of an 8-qubit, 2-layer QNN on the reduced MNIST-10 dataset with $\varepsilon=0.5$. QUID significantly degrades model performance on $\mathcal{D}_{test}$, even when trained on larger datasets.
@@
Applications: QUID is not only effective against indiscriminate poisoning attacks but can also be adapted for targeted attacks, where the adversary selectively flips labels of a specific class. This tactic allows the QNN to perform well on most test data while deliberately misclassifying the targeted class, serving the adversary’s objective. Additionally, QUID can also be used for secure dataset releases, as demonstrated in [30]. By injecting poisoned samples with labels determined by QUID, legitimate users can prevent unauthorized usage. Attackers unknowingly training on such datasets would produce suboptimal models, rendering their efforts ineffective.
@@
Limitations: The primary limitation of QUID lies in its reliance on the density matrix to assign optimal labels. While density matrices can be reconstructed on real hardware using techniques such as quantum state tomography, the measurement cost scales exponentially with the number of qubits. Recent studies [31] have proposed more efficient simulation techniques, but these approaches remain substantially more expensive compared to generic expectation-value-based measurement techniques. However, the latter approach comes at the cost of major information loss, which limits the effectiveness of QUID.
VII Conclusion
@@
In recent years, there has been significant research evaluating the security of quantum computers in cloud settings. In this work, for the first time, we focus on one such threat to QML models, specifically the threat of data poisoning attacks. An adversary in the cloud can significantly manipulate the performance of QNN models by tampering with the dataset. To this end, we propose a novel label poisoning attack, QUID, based on intra-class encoder state similarity (ESS), which can degrade the performance of QML models. Our extensive experiments under noise reveal that QUID significantly outperforms random poisoning and can degrade QNNs accuracy by $>90\%$.

### v3

@@ 5.1. Experimental Setup
5.2. Analysis
Angle vs. Amplitude Encoding: To evaluate how effectively the proposed ESS (Section 4.1) can be used to measure the performance of state preparation circuits on any given dataset, we compare the intra-class state similarity for angle and amplitude encoding under varying noise levels ($p$: error probability) on a reduced MNIST-10 dataset with a latent dimension of 16. Considering that amplitude encoding can encode $2^{n}$ features in $n$ qubits, for fairness, we used a 4-qubit circuit for angle encoding, with each qubit having four rotation gates ($R_{Z}$, $R_{X}$, $R_{Z}$, $R_{X}$), to provide a better comparison against amplitude encoding. Fig. 4 shows the performance comparison of both encoding techniques. The results are interesting, as amplitude encoding performs worse under noiseless conditions. However, it is evident that amplitude encoding outperforms angle encoding when encoding larger datasets, likely due to the loss of information caused by applying a high number of sequential rotation gates in angle encoding which is further amplified by noise. Conversely, angle encoding performs well for smaller datasets, as demonstrated in Fig. 3, where two features are encoded per qubit.
Label-flip vs. Bi-level: Even though QUID is primarily a label-flipping (LF) technique, we compare its performance against random bi-level poisoning. In this method, apart from flipping labels, we also replace samples of the poisoned data ($\mathcal{D}_{p}$) i.e. $\{x_{i}\mid x_{i}\in\mathcal{D}_{p}\subset\mathcal{D}_{tr}\}$ ($\varepsilon=0.5$), with random features and calculate the optimal poisoned labels for these random features. Table 3 shows the performance comparison of label-flipping versus bi-level poisoning when training a 4-qubit QNN on various datasets under noise with error rate $p=0.1$. Our experiments indicate that, in most cases, adding randomly generated poisoned samples (PS) actually leads to better performance (i.e. less accuracy & loss degradation) compared to the scenario of using only LF with QUID. For the remaining analysis, we use only the label-flipping technique.
Table 4. Test accuracies for QUID when trained on noiseless and noisy environment on different datasets with $\varepsilon=0.5$. QUID performs better than random label flipping under noisy conditions, demonstrating it’s robustness to noise.
@@ 5.2. Analysis
Table 5. Test accuracies of QUID with poison ratio $\varepsilon=0.5$ on real hardware noise model.
@@
Noiseless vs. Noisy Simulations: Table 4 presents the performance comparison of QUID when training a QNN under both noiseless and noisy conditions. From the table, it is evident that QUID demonstrates significantly better performance under noisy conditions compared to noiseless conditions. Specifically, under noise, QUID leads to an additional accuracy degradation of up to 24%, whereas random flipping causes only about 2% degradation. This indicates that even under noise, QUID remains effective and demonstrates a significant advantage over classical approaches.
Real Hardware Noise: Given the limited availability of real quantum devices, we directly integrated noise models from actual quantum hardware. Specifically, we employed the noise models of IBM_Kyiv and IBM_Brisbane, and incorporated them into the qiskit.aer device to simulate the noise characteristics of these devices. Table 5 lists the performance of QUID and random label-flipping on real hardware noises. We observe that even in this scenario, QUID is able to significantly degrade the performance of the baseline model by $\approx 73\%$ on IBM_Brisbane’s noise and $\approx 72\%$ on IBM_Kyiv’s noise.
Table 6. Comparison of test accuracy after training QNNs with various PQC architectures on different poison ratios ($\varepsilon$) under noise with error rate $p=0.05$. Cplx. refers to the complexity of the PQC in terms of circuit gate count and depth.
@@
Poison Ratio ($\boldsymbol{\varepsilon}$): We also evaluate the performance of QUID using different poison ratios, which helps determine the threshold beyond which the model’s performance decreases drastically. All training in this evaluation is conducted under noise ($p=0.05$) using the SPSA gradient calculation method. Table 6 lists the test accuracies of various QNNs trained across different datasets and poison ratios ($\varepsilon$). The main reason for selecting PQC-1, PQC-6, and PQC-8 is to demonstrate how differences in expressivity and noise across various PQC architectures impact the effectiveness of QUID. PQC-1 is the least expressive, while PQC-6 is the most expressive (Sim et al., 2019), but with higher gate count and depth, leading to greater noise accumulation. As shown in Table 6, this results in QUID being less effective in PQC-6, especially for higher epsilon values, as noise eventually randomizes the outputs. Conversely, for PQC-1, the QUID technique remains highly effective, with accuracy dropping to sub-1%. Consequently, we selected PQC-1 for noisy training on the other datasets.
Another important observation is that with a small poison ratio ($\varepsilon\leq 0.1$), there appears to be minimal performance degradation. However, for $\varepsilon>0.5$, the performance suffers diminishing returns, as the trained model on QUID achieves a test accuracy less than $20\%$, worse than random prediction ($25\%$ for 4-class dataset), rendering the model completely unusable for the victim. In all cases, the proposed QUID technique performs significantly better than the random label-flipping technique.
Table 7. Test accuracies for QUID when trained w/ and w/o SS-DPA (Levine and Feizi, 2021), a SOTA defense against label-flipping attacks.
@@
Dataset Baseline w/o Def. SS-DPA w/o Def. SS-DPA
@@
Defense Evaluation: Although label-flipping attacks are relatively easy to defend against in the classical domain using data sanitization techniques (Taheri et al., 2020; Steinhardt et al., 2017), implementing such defenses in the current quantum cloud environment poses significant challenges, as the victim cannot intervene once the quantum circuit and training data ($\mathcal{D}_{tr}$) are submitted for training. Therefore, we evaluate the effectiveness of QUID against a practical training-based defense technique, SS-DPA (Levine and Feizi, 2021), which is specifically designed to mitigate label-flipping attacks. SS-DPA leverages subset aggregation and ensemble learning together with semi-supervised learning. Table 7 illustrates the performance of QUID with and without SS-DPA, trained in a noiseless setting. The hyperparameter $k$ in SS-DPA is set to 3 due to the smaller dataset sizes. While the application of SS-DPA improves performance by a small amount in most cases, it is not very effective for large $\varepsilon$. Moreover, it is extremely expensive in terms of computation overhead, as it requires the training and retraining of multiple QNNs. Future research should focus on the development of effective and efficient defense techniques tailored specifically to QMLs.
6. Discussions
@@
Scalability: To demonstrate the effectiveness of QUID for larger QNNs trained on larger datasets, we designed an 8-qubit QNN with two layers of PQC-1 (Sim et al., 2019). We trained this model on the MNIST-10 dataset, reduced to a latent dimension of $d=16$, using 2000 samples with a 70:30 train-test split. Fig. 5 shows the test loss and test accuracy of the victim model when trained on a poisoned dataset with a poison ratio $\varepsilon=0.5$, using both random label flipping and QUID’s label poisoning technique. The results clearly illustrate that QUID outperforms random label flipping techniques, even for in this scenario. Specifically, QUID results in additional $\approx 41\%$ accuracy degradation and incurs $\approx 0.4$ higher test loss compared to random flipping, further validating its effectiveness.
@@
Figure 5. Test performance of an 8-qubit, 2-layer QNN on the reduced MNIST-10 dataset with $\varepsilon=0.5$. QUID significantly degrades model performance on $\mathcal{D}_{test}$, even when trained on larger datasets.
@@
Applications: QUID is not only effective against indiscriminate poisoning attacks but can also be adapted for targeted attacks, where the adversary selectively flips labels of a specific class. This tactic allows the QNN to perform well on most test data while deliberately misclassifying the targeted class, serving the adversary’s objective. Additionally, QUID can also be used for secure dataset releases, as demonstrated in (Fowl et al., 2021). By injecting poisoned samples with labels determined by QUID, legitimate users can prevent unauthorized usage. Attackers unknowingly training on such datasets would produce suboptimal models, rendering their efforts ineffective.
@@
Limitations: The primary limitation of QUID lies in its reliance on the density matrix to assign optimal labels. While density matrices can be reconstructed on real hardware using techniques such as quantum state tomography, the measurement cost scales exponentially with the number of qubits. However, recent works (Ahmed et al., 2021; Lange et al., 2023) show that classical models can now approximate these matrices accurately. This suggests that QUID can remain practical and scalable even when applied to larger QML models. Other relevant works (Chen et al., 2021) have also proposed more efficient simulation techniques. Although expectation-value-based measurement techniques are computationally cheaper, they come at the cost of significant information loss, which limits the effectiveness of QUID.
7. Conclusion
@@
In recent years, there has been significant research evaluating the security of quantum computers in cloud settings. In this work, for the first time, we focus on one such threat to QML models, specifically, data poisoning attacks. An adversary in the cloud can significantly manipulate the performance of QNN models by tampering with the dataset. To this end, we propose a novel label poisoning attack, QUID, based on intra-class encoder state similarity (ESS), which can degrade the performance of QML models. Our extensive experiments under noise reveal that QUID significantly outperforms random poisoning, degrading QNN accuracy by more than $90\%$ and by up to $50\%$ even when classical defense techniques are applied. Thus, future work should focus on developing robust defense techniques to effectively mitigate this threat in the context of QML.